# Interaction of the Primitive Equations with Sea Ice Dynamics

### v3

@@ Abstract.
This article establishes local strong well-posedness and global strong well-posedness close to constant equilibria of a model coupling the primitive equations of ocean and atmospheric dynamics with Hibler’s viscous-plastic sea ice model. In order to treat the coupling conditions, an approach involving the hydrostatic Dirichlet and Dirichlet-to-Neumann operator is developed. Mapping properties of the latter operators are investigated for the first time and are of central importance for showing that the operator associated with the linearized coupled system admits a bounded $\mathcal{H}^{\infty}$-calculus on suitable $\mathrm{L}^{q}$-spaces. Quasilinear methods allow then to obtain the strong well-posedeness results described above.
@@ 1. Introduction
The mathematical analysis of the primitive equations of the large-scale atmosphere and ocean dynamics was pioneered by Lions, Temam and Wang in [32, 33]. Moreover, Lions, Temam and Wang introduced and analyzed a coupled atmosphere-ocean model with wind-driven boundary conditions [34, 35]. Ever since then, the study of coupled atmosphere-ocean models has been an active field of research. We refer here for instance to the handbook article of Temam and Ziane [47]. In this article, for the first time, we rigorously investigate the interaction problem of the atmosphere and the ocean when additionally considering sea ice as a thin layer coupled to the atmosphere and the ocean.
@@
In a seminal paper, Cao and Titi [10] established global strong well-posedness of the primitive equations for initial data in $\mathrm{H}^{1}$ by means of energy methods. For related results, we refer to the work of Kukavica and Ziane [29]. A different approach to the primitive equations by means of evolution equations was developed by Hieber and Kashiwabara [26] as well as Giga, Gries, Hieber, Hussein and Kashiwabara [17, 18, 19]. Details of this approach and further references can be found in [24]. We also refer to the article [31] for a survey of results and further references.
In 1979, Hibler [22] proposed the governing equations of large-scale sea ice dynamics. They were investigated numerically by different communities, see e. g. [46, 37, 38, 45]. For surveys on the modeling of sea ice, we refer to [44] and [20]. We also refer here to the recent article of Piersanti and Temam [40] on the modeling and analysis of the dynamics of grounded shallow ice sheets.
This article is concerned with the analysis of a coupled atmosphere-sea ice-ocean model. More precisely, the primitive equations for the ocean and atmosphere dynamics are coupled with the sea ice equations as introduced in [22]. For a precise formulation of the coupling conditions, see (3.1), (3.2) and (3.3) in Section 3. For related results on the primitive equations subject to stochastic wind driven boundary conditions, we refer to the recent work [7].
It is the aim of this article to show that the coupled system of the ocean, the atmosphere and sea ice, see (3.4) below, has a unique, local, strong solution, and that it is globally strongly well-posed for initial data close to constant equilibria.
To this end, we rewrite the system of equations as a quasilinear evolution equation subject to coupling conditions. In order to treat these coupling conditions, we develop an approach involving the hydrostatic Dirichlet and Dirichlet-to-Neumann operator. These operators have been studied intensively in the parabolic situation and also for fluids but not for geophysical flows as the primitive equations. The results on well-posedness of the stationary hydrostatic Stokes problem with inhomogeneous boundary conditions developed in Section 7 will then be of central importance for our approach. Indeed, the hydrostatic Dirichlet and Dirichlet-to-Neumann operators allow to show that the operator associated with the linearized coupled system admits a bounded $\mathcal{H}^{\infty}$-calculus on suitable $\mathrm{L}^{q}$-spaces. The latter property allows to use modern quasilinear methods for solving the coupled system within the strong setting.
With regard to the global strong well-posedness close to constant equilibria, we make use of the so-called generalized principle of linearized stability, see [42] and [41]. We show that constant equilibria are normally stable by means of energy estimates and absorption arguments. More precisely, energy estimates yield that the spectrum of the total linearization $\mathcal{A}_{0}$ besides zero is contained in the left half-plane, that the set of equilibria close to constants is a $\mathrm{C}^{1}$-manifold and that zero is a semi-simple eigenvalue of $\mathcal{A}_{0}$. We also show that the spectrum of $\mathcal{A}_{0}$ is independent of $q$ and that thus the above properties of $\mathcal{A}_{0}$ obtained via energy estimates within the $\mathrm{L}^{2}$-setting transfer to the $\mathrm{L}^{q}$-setting for $q\in(1,\infty)$. Note that due to the quasilinear nature of sea ice dynamics, this range of $q$ is important.
For further information concerning quasilinear evolution equations, maximal $\mathrm{L}^{p}$-regularity, bounded $\mathcal{H}^{\infty}$-calculus and decoupling techniques, we refer e. g. to [4], [13], and [41].
@@
Let us note that the rigorous analysis of the single sea ice component in our model began only recently by the article of Brandt, Disser, Haller-Dintelmann and Hieber [9] as well as the article of Liu, Thomas and Titi [36]. The associated system of equations is coupled, degenerate, quasilinear and parabolic-hyperbolic. The interaction problem of sea ice with a rigid body has been studied in [5].
Models describing the coupling of sea ice with certain flows were introduced within the geophysical community e. g. by Hibler and Bryan [23] or by Timmermann, Beckmann and Hellmer [48]. Let us emphasize that these works are concerned with the development of models and not with its analysis. For analysis of related models, we refer to [14]. Coupled models of the atmosphere and the ocean call for an investigation of boundary layers. This is left to future study, and we refer to the work of Dalibard and Gérard-Varet [11] or Dalibard and Saint-Raymond [12] for boundary layer theory in the context of geophysical flows.
@@ 2. Preliminaries
We use $x$ and $y$ as well as $x_{\mathrm{H}}:=(x,y)$ to denote horizontal coordinates, whereas $z$ is employed for the vertical coordinate and $t$ is the time variable. The sub- or superscripts $o^{\mathrm{atm}}$, $o^{\mathrm{ocn}}$ and $o^{\mathrm{ice}}$ are used to indicate whether objects $o$ in the context of the atmosphere, the ocean or sea ice are considered. For instance, ${u}_{\mathrm{atm}}$, ${v}_{\mathrm{atm}}$ and ${w}_{\mathrm{atm}}$ represent the full, horizontal and vertical velocity of the atmospheric wind, ${u}_{\mathrm{ocn}}$, ${v}_{\mathrm{ocn}}$ and ${w}_{\mathrm{ocn}}$ denote the full, horizontal and vertical velocity of the ocean, and ${v}_{\mathrm{ice}}$ is the horizontal velocity of the sea ice. With $v=({v}_{\mathrm{ice}},h,a)$, the principle variable of the system will be denoted by $u=({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},v)=({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)$, while the respective pressure terms for the atmosphere and the ocean will be denoted by ${\pi}_{\mathrm{atm}}$ and ${\pi}_{\mathrm{ocn}}$. For the sea ice part, with $G:=(0,1)\times(0,1)$, we also introduce the mean ice thickness $h\colon G\times[0,T]\to(\kappa_{1},\kappa_{2})$ and the mean ice compactness $a\colon G\times[0,T]\to(0,1)$, defined as the ratio of thick ice per area. We remark that $\kappa_{1}>0$ sufficiently small is a parameter indicating the transition to open water, i. e., a value of $h(x,y,t)$ less than $\kappa_{1}$ means that at $(x,y)\in G$ at time $t$ there is open water. In contrast, $\kappa_{2}>0$ sufficiently large denotes an upper bound for the mean ice thickness.
@@
For $\Omega=G\times(a,b)$, $-\infty<a<b<\infty$, we denote by $\overline{v}$ the vertical average of $v$, i. e., $\overline{v}:=\frac{1}{b-a}\int_{a}^{b}v(\cdot,\cdot,\xi)\,\mathrm{d}\xi$. The primitive equations are modeled on the hydrostatic solenoidal $\mathrm{L}^{q}$-spaces given by the $\mathrm{L}^{q}$-closure of the smooth hydrostatic functions, so following the approach developed in [26, Sections 3 and 4], we set
@@
and horizontal periodicity is incorporated by the function spaces $\mathrm{C}_{\mathrm{per}}^{\infty}(\overline{\Omega})$ and $\mathrm{C}_{\mathrm{per}}^{\infty}(\overline{G})$, see also [24, Section 1.5.2] for further details. In that respect, for $p,q\in(1,\infty)$ and $s\in[0,\infty)$, we define the Bessel potential spaces with horizontal periodicity by
@@ 3. Main Results for the Coupled System
This is in accordance with a plane Couette flow for a Newtonian fluid, where the stress tensor is given by $\mathrm{T}-\pi\mathrm{I}_{3}$, with $\mathrm{T}$ denoting the usual Cauchy stress tensor, compare e. g. [43, Section 7.2]. The condition on the equality of the velocities on the interface of the ocean and the sea ice can be expressed by
@@
Before providing the complete coupled system, we briefly discuss the internal ice stress. We follow [22] and refer e. g. to [9, Section 2] for a more thorough treatment. In fact, the viscous-plastic rheology is expressed by a constitutive law linking the internal ice stress $\sigma$ and the deformation tensor $\varepsilon$ via an internal ice strength $P$ as well as nonlinear bulk and shear viscosities such that the principal components of the stress lie on an elliptical yield curve. With $e>1$ denoting the ratio of major to minor axes, the constitutive law is given by
@@
Even though the above law represents an idealized viscous-plastic material, the viscosities become singular for $\triangle$ tending to 00. Following [9, 37], for $\delta>0$, we take the regularization $\triangle_{\delta}(\varepsilon):=\sqrt{\delta+\triangle^{2}(\varepsilon)}$ into account and define the regularized ice stress by
@@
(3.4) $\partial_{t}{v}_{\mathrm{atm}}-\Delta{v}_{\mathrm{atm}}+({{u}_{\mathrm{atm}}}\cdot\nabla){{v}_{\mathrm{atm}}}+\nabla_{\mathrm{H}}{\pi}_{\mathrm{atm}}&={f}_{\mathrm{atm}},&&\text{ on }{{\Omega}_{\mathrm{atm}}\times[0,T]},\\ \partial_{z}{\pi}_{\mathrm{atm}}&=0,&&\text{ on }{{\Omega}_{\mathrm{atm}}\times[0,T]},\\ \mathrm{div}\,{u}_{\mathrm{atm}}&=0,&&\text{ on }{{\Omega}_{\mathrm{atm}}\times[0,T]},\\ \partial_{t}{v}_{\mathrm{ocn}}-\Delta{v}_{\mathrm{ocn}}+({{u}_{\mathrm{ocn}}}\cdot\nabla){{v}_{\mathrm{ocn}}}+\nabla_{\mathrm{H}}{\pi}_{\mathrm{ocn}}&={f}_{\mathrm{ocn}},&&\text{ on }{{\Omega}_{\mathrm{ocn}}\times[0,T]},\\ \partial_{z}{\pi}_{\mathrm{ocn}}&=0,&&\text{ on }{{\Omega}_{\mathrm{ocn}}\times[0,T]},\\ \mathrm{div}\,{u}_{\mathrm{ocn}}&=0,&&\text{ on }{{\Omega}_{\mathrm{ocn}}\times[0,T]},\\ \partial_{t}{v}_{\mathrm{ice}}-\frac{1}{m}\mathrm{div}_{\mathrm{H}}\,\sigma_{\delta}+({{v}_{\mathrm{ice}}}\cdot\nabla_{\mathrm{H}}){{v}_{\mathrm{ice}}}&=\frac{1}{m}{\tau}_{\mathrm{atm}}+\frac{1}{m}{\tau}_{\mathrm{ocn}}-g\nabla_{\mathrm{H}}H,&&\text{ on }{G\times[0,T]},\\ \partial_{t}h-d_{\mathrm{h}}\Delta_{\mathrm{H}}h+\mathrm{div}_{\mathrm{H}}\,({v}_{\mathrm{ice}}h)&=\mathrm{S}_{\mathrm{h}},&&\text{ on }{G\times[0,T]},\\ \partial_{t}a-d_{\mathrm{a}}\Delta_{\mathrm{H}}a+\mathrm{div}_{\mathrm{H}}\,({v}_{\mathrm{ice}}a)&=\mathrm{S}_{\mathrm{a}},&&\text{ on }{G\times[0,T]},\\ \mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}&={v}_{\mathrm{ice}},&&\text{ on }{G\times[0,T]}.\end{aligned}\right.$
@@
(3.5) $\partial_{\nu_{\Gamma_{u}}}{v}_{\mathrm{atm}}$ $=0\text{ on }{\Gamma_{u}},\quad\partial_{\nu_{\Gamma_{i}}}{v}_{\mathrm{atm}}=0\text{ on }{\Gamma_{i}},\quad\mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}=0\text{ on }{\Gamma_{b}},\text{ as well as}$
(3.6) $\mathrm{tr}_{\Gamma_{u}}{w}_{\mathrm{atm}}$ $=0\text{ on }{\Gamma_{u}},\quad\mathrm{tr}_{\Gamma_{i}}{w}_{\mathrm{atm}}=0\text{ on }{\Gamma_{i}},\quad\mathrm{tr}_{\Gamma_{o}}{w}_{\mathrm{ocn}}=0\text{ on }{\Gamma_{o}}\quad\text{and}\quad\mathrm{tr}_{\Gamma_{b}}{w}_{\mathrm{ocn}}=0\text{ on }{\Gamma_{b}}.$
@@
[…]
@@
${v}_{\mathrm{atm}}\in\mathrm{B}_{qp,\mathrm{per}}^{2-\nicefrac{{2}}{{p}}}({\Omega}_{\mathrm{atm}})^{2}\cap\mathrm{L}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}}),\enspace{v}_{\mathrm{ocn}}\in\mathrm{B}_{qp,\mathrm{per}}^{2-\nicefrac{{2}}{{p}}}({\Omega}_{\mathrm{ocn}})^{2}\cap\mathrm{L}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\text{ and }({v}_{\mathrm{ice}},h,a)\in\mathrm{B}_{qp,\mathrm{per}}^{2-\nicefrac{{2}}{{p}}}(G)^{4},\text{ with}$
$\partial_{\nu_{\Gamma_{u}}}{v}_{\mathrm{atm}}=\partial_{\nu_{\Gamma_{i}}}{v}_{\mathrm{atm}}=0,\enspace\mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}=0\text{ and }\mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}={v}_{\mathrm{ice}}.$
@@
[…]
(a) The statements of Theorem 3.1 and Theorem 3.2 remain valid when including Coriolis terms as in [9, Section 2]. They are omitted here for simplicity of the presentation.

### v4

# Interaction of Geophysical Flows with Sea Ice Dynamics

@@ Abstract.
This article establishes local strong well-posedness and global strong well-posedness close to constant equilibria of a model coupling the primitive equations of ocean and atmosphere dynamics with Hibler’s viscous-plastic sea ice model. In order to treat the coupling conditions, an approach involving the hydrostatic Dirichlet and Dirichlet-to-Neumann operator is developed. Mapping properties of the latter operators are investigated for the first time and are of central importance for showing that the operator associated with the linearized coupled system admits a bounded $\mathcal{H}^{\infty}$-calculus on suitable $\mathrm{L}^{q}$-spaces. Quasilinear methods allow then to obtain the strong well-posedeness results described above.
@@ 1. Introduction
In a series of articles, Lions, Teman and Wang developed a coupled atmosphere-ocean model, the so-called CAO-model, describing the atmosphere and the ocean by the primitive equations coupled by nonlinear boundary conditions, see [28, 29, 30, 31]. Their work can be regarded as the starting point of the mathematical analysis of a large class of geophysical flow models. In fact, Lions, Temam and Wang analyzed the CAO-model consisting of two primitive equations coupled by wind driven boundary conditions and proved the existence of weak solutions to this model.
@@
Starting from there, many authors studied the equations describing in particular the dynamics of the ocean by the incompressible primitive equations. In a seminal paper, Cao and Titi [10] established global strong well-posedness of the incompressible primitive equations for initial data in $\mathrm{H}^{1}$ by energy methods. For related results, we refer to the work of Kukavica and Ziane [26]. A different approach to the primitive equations by means of evolution equations was developed by Hieber and Kashiwabara [25] as well as Giga, Gries, Hieber, Hussein and Kashiwabara [18, 19, 20].
The primitive equations subject to stochastic wind driven boundary conditions have been investigated in [5], adapting an approach by Da Prato and Zabcyk [14] for stochastic boundary conditions to the situation of geophysical flows. More precisely, consider the primitive equations, where the standard deterministic boundary conditions are replaced by a stochastic boundary condition modeling the wind as $\partial_{z}V=h_{b}\partial_{t}\omega$, with a function $h_{b}$ defined at the boundary and $\omega(t)=\sum_{n=1}^{\infty}\langle g,e_{n}\rangle W(t)e_{n}$ for a suitable function $g$ and a cylindrical Wiener process $W$ on a separable Hilbert space $H$ with orthonormal basis $(e_{n})$. Following Da Prato and Zabcyk [14], one can express a solution $V$ to the primitive equations subject to the above stochastic boundary condition by a solution to the equation
$\mathrm{d}Z(t)+AZ(t)\,\mathrm{d}t=A[Nh_{b}(t)g]\,\mathrm{d}W(t),\quad Z(0)=Z_{0},$
subject to deterministic boundary conditions. Here $N$ is the so-called Neumann operator mapping deterministic inhomogeneous boundary data to the solution of the associated stationary hydrostatic Stokes problem, and $A$ denotes the hydrostatic Stokes operator. A similar method has been used in [23].
In this article, we turn our attention to an extension of the above CAO-model. Indeed, we include sea ice dynamics into the CAO-model and consider a coupled atmosphere-sea ice-ocean model. We then analyze this new coupled model, called the CIAO-model, and establish local strong well-posedness for large data and global strong well-posedness for data close to constant equilibria.
@@
Large scale sea ice dynamics is very often described by using Hibler’s model [22]. There the dynamics of the velocity, the thickness and the concentration of sea ice is described by a set of quasilinear equations. They were investigated numerically by different communities, see e. g. [33, 37]. For a survey on the modeling of sea ice, we refer to [21]. Let us note that the rigorous analysis of the Hibler’s sea ice model began only recently by the article of Brandt, Disser, Haller-Dintelmann and Hieber [8], see also [7], as well as the article of Liu, Thomas and Titi [32]. The associated system of equations is coupled, degenerate, quasilinear and parabolic-hyperbolic. The interaction problem of sea ice with a rigid body has been studied in [4].
In the present CIAO-model, the primitive equations for the ocean and the atmosphere are coupled with Hibler’s quasilinear equations for sea ice. We investigate the interaction problem of the atmosphere and the ocean when additionally considering sea ice as a thin layer coupled to the atmosphere and the ocean. For a precise formulation of the coupling conditions, see (3.1), (3.2) and (3.3) in Section 3.
Our approach to the CIAO-model can be described as follows: Motivated by the quasilinear structure of Hibler’s model, we write the CIAO-model as a quasilinear evolution equation, subject to coupling conditions. In order to treat these coupling conditions, inspired by the study of the Neumann operator $N$ as in [5], we develop a decoupling approach involving the hydrostatic Dirichlet and Dirichlet-to-Neumann operator. Such operators have been studied intensively in the parabolic situation and also for fluids but not for geophysical flows as the primitive equations. Our well-posedness results of the stationary hydrostatic Stokes problem with inhomogeneous boundary conditions shown in Section 7 are of central importance. Indeed, the hydrostatic Dirichlet and Dirichlet-to-Neumann operators allow to show that the linearized operator associated with the coupled system admits a bounded $\mathcal{H}^{\infty}$-calculus on suitable $\mathrm{L}^{q}$-spaces. This enables us to use modern quasilinear methods for solving the coupled system in the strong setting.
At this point, some comments on the background of these quasilinear methods are in order. The so-called Da Prato-Grisvard theorem [12] on the sum of two operators is a classical milestone in the theory of maximal regularity of solutions of evolution equations. It reads as follows: Let $A$ be the generator of a bounded analytic semigroup $(\mathrm{e}^{tA})_{t\geq 0}$ on a Banach space $X$ with domain $D(A)$ such that $0\in\rho(A)$, and consider $\theta\in(0,1)$ as well as $1\leq q<\infty$. Then there exists a constant $C>0$ such that for all $f\in\mathrm{L}^{q}(\mathbb{R}_{+};D_{A}(\theta,q))$, the function $u$ given for $t>0$ by $u(t):=\int_{0}^{t}\mathrm{e}^{(t-s)A}f(s)\,\mathrm{d}s$ satisfies the estimate
$\|Au\|_{\mathrm{L}^{q}(\mathbb{R}_{+};D_{A}(\theta,q))}\leq C\cdot\|f\|_{\mathrm{L}^{q}(\mathbb{R}_{+};D_{A}(\theta,q))}.$
Here $D_{A}(\theta,q)$ is defined by $D_{A}(\theta,q)=\{x\in X:[x]_{\theta,q}:=(\int_{0}^{\infty}\|t^{1-\theta}A\mathrm{e}^{tA}x\|_{X}^{q}\nicefrac{{\,\mathrm{d}t}}{{t}})^{1/q}<\infty\}$. When equipped with the norm $\|x\|_{\theta,q}=\|x\|_{X}+[x]_{\theta,q}$, the space $D_{A}(\theta,q)$ becomes a Banach space and coincides with the real interpolation space $(X,D(A))_{\theta,q}$. In addition, if $A$ is invertible, then the norm of the real interpolation space is equivalent to the homogeneous norm $\cdot]_{\theta,q}$. This result acted then as a starting point for the very many developments concerning the maximal regularity approach to evolution equations. We refer here e. g. to the work of Da Prato and Grisvard [13] on abstract nonlinear evolution equations.
Motivated free boundary problems in fluid dynamics, a homogeneous version of the Da Prato-Grisvard theorem was developed in [11]. More precisely, if $D_{A}(\theta,q)$ is endowed with the homogeneous norm $\cdot]_{\theta,q}$, then a global maximal $L^{q}$-regularity result was deduced for $1\leq q<\infty$ without assuming invertibility of the generator of the associated semigroup. This yields in particular a framework allowing for global-in-time control of the change of Eulerian to Lagrangian coordinates, which is decisive in various free boundary problems in fluid dynamics, see [11] for more information. In particular, the case $q=1$ is of crucial importance for global existence results for certain free boundary problems in the critical space $\mathrm{L}^{1}(\mathbb{R}_{+};\dot{\mathrm{B}}^{s}_{p,1}(\mathbb{R}_{+}))$.
Further important applications of the classical Da Prato-Grisvard theorem [12] concern time periodic solutions to quasilinear evolution equations. In [9], the Da Prato-Grisvard theorem was extended to the quasilinear setting to show the existence of a unique time periodic strong solution to quasilinear evolution equations in $D_{A(0)}(\theta,q)$ provided the underlying linearized operator $A(0)$ generates a bounded analytic semigroup on $X$ with $0\in\rho(A(0))$, the nonlinearities satisfy certain regularity and Lipschitz conditions, and the periodic forcing terms are sufficiently small. For further applications, see also [24].
@@ 2. Preliminaries
We use $x$ and $y$ as well as $x_{\mathrm{H}}:=(x,y)$ to denote horizontal coordinates, whereas $z$ is employed for the vertical coordinate and $t$ is the time variable. The sub- or superscripts $^{\mathrm{atm}}, oo$$^{\mathrm{ocn}}$ and $o^{\mathrm{ice}}$ are used to indicate whether objects $o$ in the context of the atmosphere, the ocean or sea ice are considered. For instance, ${u}_{\mathrm{atm}}$, ${v}_{\mathrm{atm}}$ and ${w}_{\mathrm{atm}}$ represent the full, horizontal and vertical velocity of the atmospheric wind, ${u}_{\mathrm{ocn}}$, ${v}_{\mathrm{ocn}}$ and ${w}_{\mathrm{ocn}}$ denote the full, horizontal and vertical velocity of the ocean, and ${v}_{\mathrm{ice}}$ is the horizontal velocity of the sea ice. With $v=({v}_{\mathrm{ice}},h,a)$, the principle variable of the system will be denoted by $u=({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},v)=({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)$, while the respective pressure terms for the atmosphere and the ocean will be denoted by ${\pi}_{\mathrm{atm}}$ and ${\pi}_{\mathrm{ocn}}$. For the sea ice part, with $G:=(0,1)\times(0,1)$, we also introduce the mean ice thickness $h\colon G\times[0,T]\to(\kappa_{1},\kappa_{2})$ and the mean ice compactness $a\colon G\times[0,T]\to(0,1)$, defined as the ratio of thick ice per area. We remark that $\kappa_{1}>0$ sufficiently small is a parameter indicating the transition to open water, i. e., a value of $h(x,y,t)$ less than $\kappa_{1}$ means that at $(x,y)\in G$ at time $t$ there is open water. In contrast, $\kappa_{2}>0$ sufficiently large denotes an upper bound for the mean ice thickness.
@@
For $\Omega=G\times(a,b)$, $\infty<a<b<\infty$, we denote by $\overline{v}$ the vertical average of $v$, i. e., $\overline{v}:=\frac{1}{b-a}\int_{a}^{b}v(\cdot,\cdot,\xi)\,\mathrm{d}\xi$. The primitive equations are modeled on the hydrostatic solenoidal $\mathrm{L}^{q}$-spaces given by the $\mathrm{L}^{q}$-closure of the smooth hydrostatic functions, so following the approach developed in [25, Sections 3 and 4], we set
@@
and horizontal periodicity is incorporated by the function spaces $\mathrm{C}_{\mathrm{per}}^{\infty}(\overline{\Omega})$ and $\mathrm{C}_{\mathrm{per}}^{\infty}(\overline{G})$. In that respect, for $p,q\in(1,\infty)$ and $s\in[0,\infty)$, we define the Bessel potential spaces with horizontal periodicity by
@@ 3. Main Results for the Coupled System
This is in accordance with a plane Couette flow for a Newtonian fluid, where the stress tensor is given by $\mathrm{T}-\pi\mathrm{I}_{3}$, with $\mathrm{T}$ denoting the usual Cauchy stress tensor, compare e. g. [36, Section 7.2]. The condition on the equality of the velocities on the interface of the ocean and the sea ice can be expressed by
@@
Concerning the coupling conditions, the form of the force due to the atmospheric wind in (3.1) is taken from Hibler’s classical article [22]. For the coupling of the sea ice and the ocean, it is assumed, as written above, that the force exerted by the ocean on the sea ice is proportional to the shear rate in (3.2), which corresponds to a balance of forces at the sea ice-ocean interface. The continuity of the sea ice and ocean velocities at the interface as expressed in (3.3) amounts to saying that sea ice and ocean move with the same velocity at the interface. This is inspired by the consideration of fluid-structure interaction problems where the structure is located at a part of the fluid boundary.
Before providing the complete coupled system, we briefly discuss the internal ice stress. We follow [22] and refer e. g. to [8, Section 2] for a more thorough treatment. In fact, the viscous-plastic rheology is expressed by a constitutive law linking the internal ice stress $\sigma$ and the deformation tensor $\varepsilon$ via an internal ice strength $P$ as well as nonlinear bulk and shear viscosities such that the principal components of the stress lie on an elliptical yield curve. With $e>1$ denoting the ratio of major to minor axes, the constitutive law is given by
@@
Even though the above law represents an idealized viscous-plastic material, the viscosities become singular for $\triangle$ tending to 0. Following [8, 33], for $\delta>0$, we take the regularization $\triangle_{\delta}(\varepsilon):=\sqrt{\delta+\triangle^{2}(\varepsilon)}$ into account and define the regularized ice stress by
@@
[…]
@@
[…]
@@
$\mathcal{X}_{1}$$\{$ $({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)\in(\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{atm}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}}))\times{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\times\mathrm{H}_{\mathrm{per}}^{{2},q}(G)^{2}\times\mathrm{H}_{\mathrm{per}}^{{2},q}(G)\times\mathrm{H}_{\mathrm{per}}^{{2},q}(G):$
@@
${v}_{\mathrm{atm}}\in\mathrm{B}_{qp,\mathrm{per}}^{2-\nicefrac{{2}}{{p}}}({\Omega}_{\mathrm{atm}})^{2}\cap\mathrm{L}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}}),\kern 5.0pt{v}_{\mathrm{ocn}}\in\mathrm{B}_{qp,\mathrm{per}}^{2-\nicefrac{{2}}{{p}}}({\Omega}_{\mathrm{ocn}})^{2}\cap\mathrm{L}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\text{ and }({v}_{\mathrm{ice}},h,a)\in\mathrm{B}_{qp,\mathrm{per}}^{2-\nicefrac{{2}}{{p}}}(G)^{4},\text{ with}$
$\partial_{\nu_{\Gamma_{u}}}{v}_{\mathrm{atm}}=\partial_{\nu_{\Gamma_{i}}}{v}_{\mathrm{atm}}=0,\kern 5.0pt\mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}=0\text{ and }\mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}={v}_{\mathrm{ice}}.$
@@
Remark 3.6.
(a) The statements of Theorem 3.2 and Theorem 3.4 remain valid when including Coriolis terms as in [8, Section 2]. They are omitted here for simplicity of the presentation.